OpenStudy (anonymous):

How do you find the linear approximation of tan(3 degrees) using differentials?

OpenStudy (freckles):

\[3 ^{\circ}=\frac{3 \cdot \pi}{180}= \frac{\pi}{60} \\ \text{ I would find the tangent line to } f(x)=\tan(x) \text{ at } x=0 \text{ since } \frac{\pi}{60} \text{ is pretty close to zero }\]

OpenStudy (anonymous):

so it would be 3-0 then?

OpenStudy (freckles):

to zero <--got cut off at that bit

OpenStudy (freckles):

3-0? what does that mean?

OpenStudy (anonymous):

the \[\Delta x\]

OpenStudy (freckles):

the tangent line to f at x=a is \[y=f'(a)(x-a)+f(a)\] which we will use as an approximation for f \[f(x) \approx f'(a)(x-a)+f(a) \text{ for values close to } x=a\]

OpenStudy (anonymous):

thank you!